### v1

# SMARTER: A Data‑efficient Framework to Improve Toxicity Detection with Explanation via Self-augmenting Large Language Models

Abstract
WARNING: This paper contains examples of offensive materials. Toxic content has become pervasive on social media platforms. We introduce SMARTER, a data-efficient 2-stage framework for explainable content moderation using Large Language Models (LLMs). In Stage 1, we leverage LLMs’ own outputs to generate synthetic explanations for both correct and incorrect labels, enabling alignment via preference optimization with minimal human supervision. In Stage 2, we refine explanation quality through cross-model training, allowing weaker models to align stylistically and semantically with stronger ones. Experiments on 3 benchmark tasks (HateXplain, Latent Hate, and Implicit Hate) demonstrate that SMARTER enables LLMs to achieve up to 13.5% macro-F1 improvement over standard few-shot baselines with a fraction of the full training data. Our framework offers a scalable strategy for low-resource settings by harnessing LLMs’ self-improving capabilities for both classification and explanation.
SMARTER: A Data‑efficient Framework to Improve Toxicity Detection with Explanation via Self-augmenting Large Language Models
Huy Nghiem, Advik Sachdeva, Hal Daumé III University of Maryland {nghiemh, asachde1, hal3}@umd.edu
1 Introduction
In recent years, social media platforms have enabled millions of users to virtually interact at a global scale. While a vital channel to disseminate information and knowledge, these platforms also facilitate the spread of harmful materials Hwang and Kim (2015); Cinelli et al. (2021); Nguyen et al. (2024). Arguably the most concerning among them is toxic content, which could induce serious negative psychological impact on the audience Castaño-Pulgarín et al. (2021); Windisch et al. (2022); Tontodimamma et al. (2021). As a result, content moderation is crucial to detect and mitigate its harm.
Toxic content encompasses a wide spectrum of terminologies whose definitions vary by platform: hate speech, cyberbully, sexist, racist etc. Gelber (2021); Anjum and Katarya (2024). Typically, human moderators must manually adjudicate the target content, an inefficient process taxing to the moderators themselves Baker et al. (2020); Spence et al. (2023). Recently, social media platforms can significantly reduce this overhead by integrating machine learning models to automate the detection pipeline d’Sa et al. (2020); Sumanth et al. (2022).
While capable of achieving impressive classification performance, these increasingly sophisticated models require ample resources to train, especially when the target label space cover nuanced concepts Fortuna et al. (2021); Sap et al. (2021); Zhao et al. (2021). Developing a robust dataset for training typically demands significant effort in curation and human annotation, often involving iterative updates to adapt the model to evolving linguistic trends Toraman et al. (2022); Bespalov et al. (2023); Rad et al. (2025). In addition, this pipeline is limited to decision only. Failure to provide users with human-understandable signals renders the adjudication process opaque, potentially inviting further issues on transparency and trustworthiness.
In this paper, we address the pressing need for intuitive explanations as a cornerstone of content moderation by leveraging Large Language Models (LLMs), which have shown remarkable capabilities in reasoning-related tasks Wei et al. (2021); Yao et al. (2024). Specifically, our contributions are:
  • $\diamond$
    We introduce SMARTER [^1] Our repository can be found at https://github.com/hnghiem-nlp/smarter_public (Self-augMentAtion Regimen Towards Efficient Content ModeRation), a 2-stage framework where (1) each LLM self-augments on few-shot data using preference optimization with synthetic explanations, (2) cross-model refinement to further improvement by training one LLM with responses generated by another on held-out sets.
  • $\diamond$
    Experiments are conducted on 3 benchmark datasets (HateXplain, Latent Hate, Implicit Hate) using open-source LLMs (Llama-3.1-8B-Instruct and COT-T5-XL), in few-shot settings and explainable outputs.
Our results show that leveraging multiple LLMs enables high-quality, explainable content moderation even with limited training data. SMARTER thus provides a scalable and practical approach for low-resource NLP practitioners to achieve both strong performance and explanability.
2 Related Works
Robust Toxicity Detection
Given the diverse definitions of toxicity, prior work has explored cross-domain robustness via data augmentation and architectural innovations Caselli et al. (2021); Sarkar et al. (2021); Pan et al. (2020); Toraman et al. (2022). Antypas and Camacho-Collados (2023) show that training on combined datasets improves generalization, while Nghiem et al. (2024) demonstrate that incorporating label definitions supports low-shot adaptation in BERT-based models.
Recent work has begun leveraging LLMs for content moderation, given their strong zero- and few-shot performance on toxic classification tasks Masud et al. (2024); Jahan et al. (2024); Roy et al. (2023); Kumarage et al. (2024). Beyond accuracy, LLMs can generate textual explanations, offering interpretable signals for moderators and users alike Di Bonaventura et al. (2024); Yang et al. (2023). Leveraging these insights, we propose a framework that trains open-source LLMs to jointly perform classification and explanation across toxic domains, enhancing transparency in moderation workflows.
Aligning LLMs to Human Preferences
Despite their impressive capabilities, LLMs still have their limitations: they may generate unhelpful responses to user queries Liu et al. (2023); Shen et al. (2023), or sycophantic outputs that pander to the user’s perceived preference when prompted Papadatos and Freedman (2024); Ranaldi and Pucci (2023) . Various techniques have been introduced to align LLMs towards preferred generative patterns: from finetuning approaches Dai et al. (2023); Sun et al. (2023); Wang et al. (2023) to interactive mechanisms (prompting) Furniturewala et al. (2024); Guo et al. (2022). A promising line of research showcases LLMs’ self-improving potential via their own generated synthetic data Tang et al. (2023); Li et al. (2023). Combining these approaches, we exploit LLMs’ sycophantic behavior to augment limited training data resource and improve downstream performance via alignment techniques.
3 Overview of Framework
As shown in Algorithm 1, our framework is organized into two stages. First, we focus on optimizing the LLMs’ classification performance through alignment tuning using self-augmenting synthetic data generation. In the second stage, we refine the quality of the generated explanations by cross-model refinement, leveraging the synergy between different LLMs. Section 4 and 5 present the experimental setup and findings for Stage 1. Section 6 focuses on cross-model improvement in Stage 2.
Algorithm 1 High-Level Overview of SMARTER
1:Input: LLMs $\mathcal{L}$, performance metric $\mathcal{M}$ Datasets $\mathcal{D}_{\text{train}}\text{ (with explanation)},\mathcal{D}_{\text{val}},\mathcal{D}_{\text{test}}$,
2:Optional: Initial SFT pretraining on relevant data
3:Stage 1: Individual LLM Self-Augmentation
4:for $K\in\{K_{\text{start}},...,K_{\text{end}}$ } do
5:  Sample $K$-shot subset $\mathcal{D}^{(K)}$
6:  for all $\text{LLM}_{i}\in\mathcal{L}$ do
7:   SFT on $\mathcal{D}^{(K)}$
8:   Collect chosen responses for gold labels
9:   Collect rejected responses for incorrect labels
10:   Align via preference optimization using synthetic preference data; evaluate on $\mathcal{D}_{\text{val}}$
11:  end for
12:end for
13:Stage 2: Cross-Model Refinement
14:for $\text{LLM}_{a}\neq\text{LLM}_{b}\in\mathcal{L}$ do
15:  Select $K_{check}\in\{K_{\text{start}},...,K_{\text{end}}\}\rightarrow\mathcal{D}^{(K_{check})}$
16:  SFT $\text{LLM}_{a}$ on $\mathcal{D}^{(K_{check})}$
17:  Self-augment $\text{LLM}_{a}$ with responses from $\text{LLM}_{b}$ on $\mathcal{D}^{(K_{check}^{\prime})}\neq\mathcal{D}^{(K_{check})}$
18:end for
19:return Model with best performance $\mathcal{M}$ on $\mathcal{D}_{val/test}$
4 Preliminary Setup
In this section, we provide details about the empirical setup on data, model, and a description of classification technique via prompting.
4.1 Data
We select 2 benchmark datasets that cover distinct yet still relevant concepts to toxic content moderation and derive from them 3 classification tasks.
  • $\diamond$
    HateXplain, introduced by Mathew et al. (2021). Consisting of posts collected from 2 social media sites Gab and Twitter (now X), this dataset contains 20,148 samples annotated by crowdworkers of the Amazon Mechanical Turks (AMT) platform on 3 labels Hate, Offensive and Normal.
  • $\diamond$
    Latent Hate, curated by ElSherief et al. (2021), focuses on covert hate speech collected from Twitter. With 19,112 samples also annotated by AMT workers, this dataset’s primary label space contains 3 high-level categories: Not Hate, Implicit Hate, Explicit Hate. The authors also assigned the subset of 4,153 tweets of the Implicit Hate category a label among 6 secondary fine-grained categories: White Grievance, Incitement to Violence, Inferiority Language, Irony, Stereotypes and Misinformation, Threatening and Intimidation.
Detailed definition of the terminologies associated with the labels are also included, allowing us to integrate these contexts into the explanation in downstream applications in a fashion similar to Yang et al. (2023) and Nghiem and Daumé (2024). We hereby refer to the 3 classification tasks based on their distinct label spaces of as HateXplain, Latent Hate, Implicit Hate respectively. Table 1 summarizes the main components of these datasets and the split for training, validation and testing. Appendix B describes these pre-processing steps.
Dataset Total Size Split Ratio K val K test Platform Label Space
HateXplain 20,148 60:20:20 50 400 G, T Normal, Offensive, Hate
Latent Hate 19,112 60:20:20 50 400 T Not Hate, Explicit Hate, Implicit Hate
Implicit Hate 4,153 50:20:30 50 150 T White Grievance, Incitement to Violence, Inferiority Language, Irony, Stereotypes and Misinformation, Threatening and Intimidation
Table 1: Datasets for evaluating our framework. Split Ratio designates the proportion of the train:validation:test split. K val and K test are the number of data points per class to perform $K$-shot sampling from the Validation and Test set. For Platform, G denotes Gab, T for X (formerly Twitter).
4.2 Models
Extending our focus on accessibility, we employ 2 open-source LLMs that have exhibited strong performance in similar classification tasks:
  • $\diamond$
    COT-T5-XL: An encoder-decoder transformer-based Vaswani (2017) LLM introduced by Kim et al. (2023). A variant of the base Flan-T5-XL Longpre et al. (2023) architecture, this model consists of 3 billion parameters and is further trained on the COT corpus, a collection of 1.8 million samples with chain-of-thought style explanations, which enhances its capabilities on a variety of reasoning tasks Kim et al. (2023).
  • $\diamond$
    Llama-3.1-8B-Instruct: A decoder-only LLM released by Meta Dubey et al. (2024). Consisting of 8 billion parameters, this model is further instruction-tuned to aligned with human preferences in addition to being trained on extensive linguistic corpora.
For brevity, we refer to the models as T5 and Llama respectively for the remaining of this paper.
4.3 Classification via Prompting
As demonstrated by Wei et al. (2022), prompting an LLM of sufficient size to provide justification before the decision enhances its downstream performance. We apply this Chain-of-Thought (COT) paradigm in our classification tasks to obtain the explanation from the LLMs necessary for content moderation. Inspired by Nghiem and Daumé (2024) and Yang et al. (2023), we use the prompt shown in Figure 1, which provides the target post, the set of labels corresponding to the task and their associated definitions, and the instruction for the LLM to provide its explanation and predicted category using this context.
### Instruction:
By considering their corresponding Definitions, label the following post with only one of these categories: categories. Provide a brief paragraph to explain step-by-step why the post should be classified with the provided Label based on the given Definitions.
If this post targets a group or entity relevant to the definition of the specified Label, explain who this target is/are and how that justifies the chosen Label.
Provide your message in the following format:
EXPLANATION: [text]
LABEL: [text]
### Definitions:
Hate: Language that targets [...]
Abusive: Language that is [...]
Normal: Does not contain [...]
### Post:
{post}
### Response: [model output]
Figure 1: Prompt template for classification tasks. The definition block should contain the set’s full label space
4.4 Auxiliary Pretraining
Leveraging existing resources, we train both T5 and Llama models on HateCOT, a dataset introduced by Nghiem and Daumé (2024), which consists of toxic posts, their labels, definitions and synthetic explanations designed to boost of-the-shelf LLMs’ performance in hate speech detection. Specifically, the models are supervised finetuned on the next token prediction tasks using inputs constructed by combining HateCOT’s data with our prompt template. Finetuning is performed by training LoRA adapters Hu et al. (2021), a parameter-efficient approach implemented by the HuggingFace library (see Appendix E), which serve as the basis for downstream experiments.
5 Stage 1: Individual Model Training
5.1 K-shot In-domain Finetuning
We experiment with different configurations of limited in-domain data. For each task, we select uniformly at random K-shot $\in\{16,32,64,128,256\}$ samples from the training set . We utilize the seed explanations by Nghiem and Daumé (2024), which was generated by OpenAI’s GPT-3.5-Turbo. Further finetuning the existing HateCOT-pretrained T5 and Llama models happens in a fashion similar to the auxiliary pretraining phase above.
Self-augmenting via Alignment Tuning
We optimize the LLMs’ performance while not having access to more training data by exploiting the models’ sycophantic tendency. More concretely, we augment the training sample size by prompting the LLMs to generate explanations conditioned on the incorrect labels (as illustrated in Figure 2). This process supplies us with synthetic post-explanation pairs that are by design dispreferred compared to the true preferred sample pairs, enabling the use of the following alignment techniques to further finetune the models.
  • $\diamond$
    Direct Preference Optimization (DPO): An offline Reinforcement Learning (RL) technique that optimizes the policy via an implicit reward model using preference data Rafailov et al. (2024). DPO requires human preference data where a pair of chosen vs. rejected responses is presented for each prompt.
  • $\diamond$
    Kahneman-Tversky Optimization (KTO): An alternative to DPO, KTO leverages prospect theory to construct human-aware losses (HALOs) to optimize the policy Ethayarajh et al. (2024). KTO only requires listwise human preference data: a data point consists of a prompt, a response and a binary flag that indicates the acceptability of the response.
Appendix E.2 offers more technical details on these techniques. Our self-augmenting methodology naturally procures training datasets in the required format for KTO by simply designating the explanation conditioned on the gold label as positive, and otherwise negative. For DPO, we match the explanations conditioned on the correct-incorrect pair sequentially to satisfy the pairwise chosen-rejected format. As an example, 100 posts of HateXplain (whose label space is of size 3), should produce 1003=300100*3=300 training samples for KTO, and 100(31)=200100*(3-1)=200 for DPO.
Figure 2: Self-augmenting pipeline: for each post, explanations are conditionally generated for the gold label and all incorrect labels using prompt template in Figure 1. For DPO, data consists of matching the explanation of the correct label with another incorrect label’s.
5.1.1 Experimental Pipeline
The experiments follow a two-stage process. First, we finetune T5 and Llama under extreme data scarcity by setting $K\in\{16,32,64,128\}$, followed by self-augmentation via DPO to compare classification performance. Then, we experiment with significantly more training data with $K=256$ and evaluate KTO alignment against DPO  [^2] Implementation for KTO trainer for T5 is not available at the time of writing, so we are only able to perform this variant for Llama. For comparison, we also finetune both LLMs with the entire available training data to obtain classification lable without explanation, denoted Full models. In addition, we also finetune ModernBERT (large) Warner et al. (2024), a decoder-only model on the full training data and obtain the classification outputs using the same test sets.
In addition, we examine 2 partial sampling strategies to circumvent the proportionate growth of self-augmenting synthetic data to label space when K is large . In the first, we construct augmented data by selecting 128 and 192 shots of samples uniformly at random from the Baseline’s 256-shot pool. In the second, instead of $K$-shot sampling, we select $K^{\prime}*|S|*(|S|-1)$ samples from all possible post-label pairs of the 256-shot Baseline pool (incorrect labels included) to create augmented data, where $K^{\prime}\in\{128,192\}$ and $|S|$ is the label space’ size. While both approaches should yield identical size of training data for the same $K^{\prime}$, the first method, denoted DPO-K, preserves post-label parity whereas the second method, denoted DPO-N, yields higher diversity of post content.
Experiments are implemented with the HuggingFace library. We sample $K=50$ shot on the held out Validation extraction of each dataset to tune hyperparamters, and evaluate on samples from the Test set as shown in Table 1. Technical details are included in Appendix E.
5.2 Classification Results
5.2.1 Results for $K<=128$
We report macro-F1–the harmonic mean between Precision and Recall–as the evaluation metric Christen et al. (2023). Figure 3 displays the bar plots of the LLM’s Baseline and DPO-enhanced performance on the test sets.
Self-augmentation with DPO generally boosts classification performance over its Baseline counterparts, with the sole reversal for T5 on Latent Hate at $K=16$ (2.1% reduction) along with static scores for T5 on HateXPlain at $K=32$, and for Llama in Latent Hate at $K=64$. The infrequency of these observations and the negligible decrease highlight our method’s consistent enhancement. In fact, increase in F1 scores go up to 14.5% for Llama (HateXplain, $K=128$) and 25% for T5 (Implicit Hate, $K=64$).
Notable is DPO self-augmentation’s ability to improve over Baselines at even $K\leq 64$, showcasing our method’s advantage in low-resource settings. Llama Baseline models appear to consistently outperform T5 across all tasks. However, with DPO self-augmentation, T5 is able to narrow the margin on Latent Hate and Implicit Hate. In contrast, Llama still dominates on HateXplain. These discrepancies suggest an interaction between dataset and model choice, and practitioners may benefit from contrasting different options with our self-augmenting techniques for best results.
Figure 3: Bar plots for $K$-shot classification experiments on 3 datasets using Llama and T5 models. Macro F1 scores and percentage change over Baseline are displayed on top. Results for Baselines and DPO-augmented variants for $K\in\{16,32,64,128\}$ are displayed on the left subfigures. Results for $K=256$ of Baseline, KTO, DPO-augmented and its other variants on various sub-sampling strategies (section 5.1.1) are shown on the right. Horizontal lines show the F1 scores for Full models that use all training data.
5.2.2 Results for $K=256$
In Figure 3, we observe the macro F1 scores for models $K=256$ on the subfigures on the right. DPO-augmentation continues to enhance Baseline models which already benefits from more training data relative to lower $K$ values, with improvement in scores reaching up to 13.2% for LLama and 14% for T5. In contrast, KTO-augmentation appears ineffectual on HateXplain and even hinders performance on Latent Hate and Implicit Hate, showing notable reduction from Baseline scores, respectively. With this, we hypothesize that DPO’s contrastive nature reinforces signals that allow the the LLM to learn from its own mistakes, thereby sharpening its reasoning ability for this type of discriminatory task. On the other hand, KTO’s reliance on implicit signals from listwise data is not sufficient to improve discernment of classes, particular when the semantic distinction between labels are nuanced as in the case of Latent Hate and Implicit Hate.
Results also show that DPO-augmented sub-sampling strategies still demonstrate generally better F1 scores than Baselines, further reinforcing the benefit of DPO-augmentation. There appears no clear winner between DPO-K and DPO-N among all 3 datasets. Augmenting with larger $K^{\prime}$ values (approaching 256) consistently yields scores closer to the maximal F1, indicating its role as a hyperparameter that practitioners need to decide for the trade-off between computational cost and benefit.
Dataset Model F1 Full F1% Data%
HateXplain Llama 0.72 89% 6%
T5 0.72 86%
ModernBERT 0.70 100% 100%
Latent Hate Llama 0.62 111% 7%
T5 0.61 107%
ModernBERT 0.61 100% 100%
Implicit Hate Llama 0.67 90% 57%
T5 0.64 92%
ModernBERT 0.64 100% 100%
Table 2: Comparison of performance and data usage of DPO-augmented vs. Full models at $K=256$. F1 Full denotes the macro F1 scores of Full models that use the entire available training data. F1% is the percentage of the F1 score attained by the corresponding DPO-augmented model over F1 Full. Data% is the percentage of size of the training data using $K=256$ shot over the full training samples. ModernBERT’s scores are presented for references.
Remarks
Our empirical results demonstrates that our DPO augmentation enables LLMs to achieve considerable performance with low resource. From Figure 3, DPO-augmented models attain high macro F1 scores at $K=64$, a very small amount of data. Table 2 further showcases that DPO-augmented models can achieve at least 86% of F1 performance (HateXplain) or even exceed that of their Full counterpart (Latent Hate) while using only fraction of the available training data at $K=256$. For practitioners, these insights suggest a resource-conservative strategy: start with low K-shot samples (64) to obtain benchmark results, then iteratively add more training samples until performance plateaus or resource is exhausted.
6 Stage 2: Cross-model Refinement
In this section, we focus on investigating methods to identify and efficiently improve the LLMs’ quality of explanations with respect to certain desired directions. We first enlist human annotators to evaluate the explanations by T5 and Llama, then experiment with further training one model with the other’s outputs to simultaneously improve performance and quality of explanations.
6.1 Human Evaluation
Due to resource constraints, we only perform this evaluation pipeline for HateXplain. We opt to collect outputs from the SFT+DPO models at $K=128$ to preserve the rest of the training data for further refinement experiments. To make a fair comparison on the outputs to agree on the predicted label, we use the prompt in Figure 7 to obtain explanations conditioned on the correct gold label. This technique allows us to obtain 342 samples with consistent labels, or 89% of the possible 1283=384128*3=384 possible samples. The other 42 samples contain discrepant predicted labels, where the LLMs revert to behaviors ingrained by previous training.
To simulate real-life deployment, we solicit 14 crowdworkers from diverse demographic backgrounds on the platform Amazon Mechanical Turks (MTurks) to annotate their preference on the explanations offered by the LLMs in this study . Annotators are encouraged to evaluate the pairs of explanation for each post based on the criteria of Clarity, Reasoning and Alignment as illustrated with the template in Figure 6. To avoid position bias, we present the pairs of explanations in randomized orders. All annotators are compensated fairly.
Label T5 Llama Row Total
Normal 25 (25.5%) 73 (74.5%) 98
Offensive 63 (49.6%) 64 (50.4%) 127
Hate 54 (46.2%) 63 (53.8%) 117
Table 3: Results of preference for explanations by MTurk annotators based on majority voting for 342 HateXplain samples.
Annotation Results
Table 3 shows the annotation results, aggregated per sample by majority voting. Across the 3 labels in HateXplain, annotators show no significant preference for explanations by either model on Offensive and Hate. On the other hand, Llama’s explanations for Normal posts are overwhelmingly preferred.
6.2 Cross-model Stylistic Refinement
Since Llama produces outputs preferred by human annotators, a natural choice is to select this model for ultimate deployment. We are nevertheless interested in examining the potential to steer one model’s explanation style to align with another model’s while maintaining existing classification performance.
Using the prompting technique described in section 6.1, we use the SFT+DPO $K=128$ variants of both Llama and T5 LLMs to generate explanations conditioned on the gold labels of the held-out 128-shot portion of the training data, which neither of these models has seen. We then perform cross-model SFT and DPO finetune using this dataset. More specifically, we further finetune the existing SFT+DPO $K=128$ T5 variant using explanations produced by its Llama counterpart, and vice versa. This step consists of both separate SFT and then DPO self-augmentation as in section 5.1.
For each test domain, we additionally train a style classifier based on the BERT architecture using explanations generated individually by T5 and Llama of the SFT+DPO variant at $K=128$ Devlin et al. (2019). A classification head is added to BERT, which leverages the [CLS] token from each explanation’s embedding to predict whether it was produced by T5 or LLaMA. Technical details are reported in Appendix E.2.
6.2.1 Cross-model Refinement Results
Figure 4: Macro-F1 scores on test portion of the 3 test sets for T5 and Llama cross-model refinement regimen. In each figure: the first and last bars are scores reported with only T5 model at $K=128$ and $K=256$ (directly from Figure 3); the middle bars are results after further finetuned using data from the complementary $K=128$ shots of counterpart model. Split-color bars: bottom color indicates the original model; top color indicates the counterpart model for additional cross-training.
Figure 4 shows the macro-F1 scores of both the original single-model SFT+DPO and cross-model finetuned variants on the test datasets.
Cross-model training enhances T5 while impairing Llama.
Applying SFT training using explanations from Llama improves performance over the baseline T5 SFT+DPO on HateXplain (F1 scores 0.62 from 0.60) and Latent Hate (0.64 from 0.62), with a slight reduction on Implicit Hate (0.54 from 0.56). On the other hand, applying SFT+DPO augmentation using Llama’s explanations from the complementary K=128 shot data allows T5 to exceed single-model performance on the full K=256 set across the board. The new maximum macro-F1 scores for T5 are 0.66 on HateXplain (up from 0.62 of single-model), 0.66 on Latent Hate (up from 0.65), and 0.61 on Implicit Hate (up from 0.59). Notably, these scores are superior to even Llama’s single-model performance at $K=256$ on HateXplain and Implicit Hate.
In contrast, SFT cross-training with T5’s output reduces LLama’s F1 scores to 0.58 from 0.63 on HateXplain while improving to 0.66 from 0.62 on Latent Hate and plateauing at 0.58 on Implicit Hate. Additional SFT+DPO cross-model augmentation slightly boosts performance on HateXplain and Implicit Hate while slightly impairing on Latent Hate. None of the cross-model Llama variants matches the single-models’ F1 scores at $K=256$.
(a) HateXplain
(b) Latent Hate
(c) Implicit Hate
Figure 5: Percentage distribution of Llama- and T5-style explanations on test sets by BERT style classifiers, pre (Base) and post cross-model (Xmod) refined.
6.2.2 Stylistic Analysis
From Figure 5, most explanations from the individually trained models exhibit their corresponding style, while cross-model SFT fine-tuning is sufficient to substantially align the outputs with the alternative’s style. With the exception of T5 on HateXplain, the application of self-augmenting DPO using cross-model data generally amplifies the stylistic presence of the base model to a small extent, suggesting that this mechanism induces a mild converging effect on the reasoning style.
Cross-model explanations improve weak categories.
We analyze how cross-model refinement enhances classification by comparing Precision, Recall, and F1-score breakdowns of the base models and their cross-trained counterparts at $K=256$ in Figure 8. Models often improve in categories where they initially struggled but their counterpart performed well, suggesting they can adopt effective reasoning patterns. However, performance may decline if the counterpart model is weak. This tendency, along with the stylistic transference shown in Figure 5, aligns with findings on the benefits of reasoning diversity by incorporating outputs from other models Hao et al. ; Chenglin et al. (2024).
7 Discussion and Conclusion
Informed by empirical insights , we offer the following recommendations for practitioners.
Cross-model refinement can further improve performance
T5’s superior F1-scores compared to individually trained models (Figure 4) suggest added value in leveraging competing models. Practitioners can train multiple LLMs using the SFT+DPO framework across data thresholds, then select a high-quality checkpoint (e.g., $K=128$ from Section 5.2.1) for further finetuning with outputs from other models. Though additional performance gain may not be guaranteed, cross-model refinement is still a venue worthy of exploration particularly when resources are limited.
Architectural diversity may offer benefits.
The differing post-training behaviors of T5 and Llama—despite being trained on the same seed explanations—may stem from their distinct architectures (encoder-decoder vs. encoder-only). We hypothesize, based on prior work, that the stronger model (Llama) enhances reasoning diversity that T5 can leverage. While further study is needed for a definitive explanation, our findings suggest practitioners may benefit from using architecturally diverse models within this framework.
Human oversight is still essential.
Beyond the initial annotation, human supervision is necessary to monitor the stylistic quality of explanations, especially in iterative-development settings where unexpected patterns may occur after multiple rounds of cross-model training.
We hope that SMARTER will enable practitioners to effectively and efficiently deploy explanation-based content moderation pipelines.
8 Limitations
We discuss several limitations in our research, along with their implications for future explorations.
Generalization beyond English
Our work focuses exclusively on English corpora. In reality, social media exists for every language. As both English-based LLMs and research literature often attract much more attention, we hope this framework will be adapted for other languages.
Model selection
While we perform extensive experiments, our selection of 2 models could be expanded. At the cost of more training resource, training more models with different architectures or generation offer extra options for performance comparison for mix-and-match cross-training. We encourage other researchers to experiment with more LLMs using our framework.
Limited scope of human validation
Due to budget constraints, we are only able to human validation for 1 dataset. In practice, it is recommended to solicit a larger pool of annotators to ensure appropriate amount of human oversight.
Risk of bias in automated explanation
While our framework aims to improve performance and versatility in content moderation, it does not alleviate the fundamental sensitive nature of this task. Content moderation could be misappropriated to suppress free speech or harm marginalized groups Dias Oliva (2020); Kozyreva et al. (2023). Other works have noted that LLMs exhibit bias behaviors in certain applications Nguyen et al. (2023); Gu et al. (2025). We urge researchers and practitioners to maintain vigilance when integrating our framework with human supervision to ensure ethic standards Lai et al. (2022); Cao et al. (2024).
9 Ethical Considerations
The authors are not aware of any ethical problems in the development of this work. This research uses publicly available dataset. We also further synonymize the content to the best of our ability to provide extra caution for privacy. We acknowledge that our framework and associated techniques could be abused for harmful purposes.
Appendix
Appendix A Amazon Turks Mechanical Annotation
We obtained IRB approval before soliciting annotation from the crowd workers. 14 workers were recruited to annotate the 342 explanations introduced in subsection 6.1. Samples are assigned at random by the AMT platform to each annotator. Due to budget constraint, we opt to collect annotation for HateXplain to preserve statistical power instead of spreading across all 3 datasets. The breakdown of the annotators’ demographic is below:
  • $\diamond$
    Gender: Female (10), Male (3), Non-binary (1)
  • $\diamond$
    Age: 18-29 (2), 30-39 (5), 40-49 (5), 50+ (2)
  • $\diamond$
    Race/Ethnicity: Asian (3), Black (2), Hispanic/Latino (1), White (8)
  • $\diamond$
    Country of residence: United States (14)
  • $\diamond$
    Education level: 2-year college or equivalent (4), 4-year college or equivalent (6), High school or equivalent (3), Master degree or above (1)
Figure 6: Template to collect annotation for preference on explanations for Amazon Mechanical Turks crowdworkers.
Appendix B Data Pre-processing
The datasets used in this work are open-source. We peruse them as research artifacts accordingly to their license. To further mitigate risks of privacy, we anonymize posts by replacing replacing user handles with the string <user> and remove URLs if they appear in the context of the posts.
We obtain the seed explanations from the repository associated with Nghiem and Daumé (2024) and merge them with the correct identifier keys.
Appendix C Obtaining Explanations for In-domain Data
The prompt template in Figure 7 is used to collect explanation conditioned on a single label. Note that more labels (and their definitions) may also be added to this section as desired.
### Instruction:
By considering the following Post, the Label and its Definition, provide a brief paragraph to explain step-by-step why the post should be classified with the provided Label based on the given Definition.
If this post targets a group or entity relevant to the definition of the specified Label, explain who this target is/are and how that justifies the chosen Label.
Provide your message in the following format:
EXPLANATION: [text]
LABEL: [text]
### Definitions:
Hate: Language that targets [...]
### Post:
{post}
### Response:
Figure 7: Prompt template to obtain explanation conditioned on a single label and its definition.
Appendix D Cross-model Analysis
(a) HateXlain
(b) Latent Hate
(c) Implicit Hate
Figure 8: Comparison of Precision, Recall, and Macro F1-scores for the T5 and Llama model families under two training regimes: self-augmentation only (Base) and cross-model refinement (+Xmod) at $K=256$. Models trained with cross-model refinement (right plot within each model family) generally show improvements in label categories where their counterparts already perform well, and conversely exhibit declines in areas corresponding to their weaknesses.
Appendix E Technical Details
E.1 Hardware
All experiments are carried out on a maximum of 2 Nvidia RTX A6000 GPUs. We download the models from their Huggingface [^3] https://huggingface.co/ checkpoints. Finetuning is implemnted with LoRA techniques Hu et al. (2021) via the TRL library  [^4] https://huggingface.co/docs/trl/en/index. The LoRA configurations for all settings are:
  • $\diamond$
    Rank: 64
  • $\diamond$
    alpha: 128
  • $\diamond$
    Dropout: 0.05
  • $\diamond$
    Target modules: $q$ and $v$ in projection layers
E.2 Preference Optimziation
Direct Preference Optimization
DPO is an alignment technique used to fine-tune large language models (LLMs) to better align with human preferences. Unlike traditional methods such as Reinforcement Learning from Human Feedback (RLHF), DPO simplifies the process by directly optimizing the policy model using a binary cross-entropy objective, thereby eliminating the need for a separate reward model and the complexities of reinforcement learning. This approach reparameterizes the reward model such that its optimal policy can be expressed in a closed form, transforming preference optimization into a classification problem Rafailov et al. (2024).
Kahneman-Tversky Optimization
KTO is another model alignment technique that aims to directly maximize the utility of a model’s generations by applying principles from Kahneman-Tversky prospect theory Kahneman and Tversky (2013). KTO evaluates outputs based on their perceived gains or losses relative to a reference point, incorporating the concept of loss aversion. This method utilizes data that simply indicates whether an output is desirable or undesirable for a given input, which can be more readily available and less expensive to collect than the paired preference data required by DPO Ethayarajh et al. (2024).
Implementation
We primarily use the implementation of the DPOTrainer and KTOTrainer classes via the TRL library hosted on the HugingFace platform. For the DPOTrainer, we set $\beta=0.1$, and the loss to be default sigmoid loss. For the KTOTrainer, we similarly set the $\beta=0.1$ and use the default KTO loss option.
E.3 Hyperparameter Tuning
For ModernBert, we use the HuggingFace Trainer and train for 4 epochs with AdamW optimizer. We use the associated tokenizer and pad the inputs to maximum length (512) for each batch.
For Base SFT finetuning, we use 3 epochs and learning rate 3e-4. We present the chosen parameters for DPO/KTO alignment tuning in Table 4, Table 5, Table 6 below.
T5 Llama
K Technique Epochs LR Epochs LR
16 DPO 3 5e-05 3 1e-05
32 DPO 3 5e-05 4 1e-05
64 DPO 3 1e-05 3 1e-05
128 DPO 3 5e-05 3 1e-05
256 DPO 4 1e-04 3 1e-05
256 KTO 3 5e-07
(a) Hyperparameters for regular training.
T5 Llama
K Technique Epochs LR Epochs LR
256 DPO-K128 3 1e-05 4 1e-04
256 DPO-K192 3 1e-04 1 7e-05
256 DPO-N128 3 5e-05 5 5e-06
256 DPO-N192 3 5e-05 5 5e-06
(b) Hyperparameters for subsampling training.
Table 4: HateXplain.
T5 Llama
K Technique Epochs LR Epochs LR
16 DPO 3 5e-05 3 1e-06
32 DPO 3 1e-04 3 1e-06
64 DPO 3 5e-05 3 1e-06
128 DPO 4 5e-05 3 1e-04
256 DPO 3 1e-04 3 1e-04
256 KTO 3 5e-07
(a) Hyperparameters for regular training.
T5 Llama
K Technique Epochs LR Epochs LR
256 DPO-K128 4 1e-04 4 1e-04
256 DPO-K192 3 1e-04 3 1e-04
256 DPO-N128 4 1e-04 3 1e-04
256 DPO-N192 4 1e-04 5 1e-04
(b) Hyperparameters for subsampling training.
Table 5: Latent Hate.
T5 Llama
K Technique Epochs LR Epochs LR
16 DPO 3 5e-05 3 5e-06
32 DPO 3 5e-05 4 5e-05
64 DPO 3 1e-04 3 1e-06
128 DPO 1 7e-05 3 1e-04
256 DPO 1 5e-05 1 5e-05
256 KTO 3 5e-07
(a) Hyperparameters for regular training.
T5 Llama
K Technique Epochs LR Epochs LR
256 DPO-K128 1 7e-05 1 1e-05
256 DPO-K192 1 7e-05 1 5e-05
256 DPO-N128 1 7e-05 1 1e-05
256 DPO-N192 1 7e-05 1 5e-05
(b) Hyperparameters for subsampling training.
Table 6: Implicit Hate
Dataset Definition
HateXplain $\diamond$ Normal: neither hate speech nor offensive. $\diamond$ Hate Speech: language which attacks, demeans, offends, threatens, or insults a group based on race, ethnic origin, religion, disability, gender, age, sexual orientation, or other traits. it is not the presence of certain words that makes the text hate speech, rather you should look the context the word is used in the text. $\diamond$ Offensive Language: usage of rude, hurtful, derogatory, obscene or insulting language to upset or embarrass people.
Latent Hate $\diamond$ Not Hate: speech or actions that do not involve any form of hatred, prejudice, or discrimination toward individuals or groups based on their characteristics. $\diamond$ Explicit Hate: openly expressed, direct forms of hatred and prejudice toward individuals or groups based on their characteristics. $\diamond$ Implicit Hate: coded or indirect language that disparages a person or group on the basis of protected characteristics like race, gender, and cultural identity.
Implicit Hate $\diamond$ White Grievance: includes frustration over a minority groups perceived privilege and casting majority groups as the real victims of racism. This language is linked to extremist behavior and support for violence. $\diamond$ Inferiority Language: implies one group or individual is inferior to another, and it can include dehumanization, denial of a person’s humanity, and toxic language that compares the target with disease, insects, animals. Related to assaults on human dignity, dominance, and declarations of superiority of the in group. $\diamond$ Irony: refers to the use of sarcasm , humor, and satire to attack or demean a protected class or individual. $\diamond$ Stereotypes And Misinformation: associate a protected class with negative attributes such as crime, or terrorism. includes misinformation that feeds stereotypes and vice versa, like holocaust denial and other forms of historical negationism. $\diamond$ Threatening And Intimidation: conveys a speaker’s commitment to a target’s pain, injury, damage, loss or violation of rights, threats related to implicit violation of rights and freedoms, removal of opportunities, and more subtle forms of intimidation.
Table 7: Definitions of toxic label categories and definitions (collected from their original publications) across datasets.
Appendix F Examples of Explanations
We provide some examples of the explanations generated during our cross-training phrase in Figure 9, Figure 10, and Figure 11.
POST:
what kind of brzen hapless *** uses an expression like "there’s no big *** there"
LABEL: HATE
(a) A sample post in HateXplain with gold label. Offensive content partially redacted.
VARIANT: T5 DPO @ K=128
EXPLANATION: The post targets the group of homosexuals by using a derogatory term "***" to insult and demean them. This language is offensive as it is rude, hurtful, and insulting. Therefore, the post is labeled as Offensive.
LABEL: Offensive
VARIANT: T5 +XMOD SFT+DPO @ K= 256
EXPLANATION: The post targets the group of individuals who identify as "brzen" and use the derogatory term "***" The language used in the post attacks and demeans this group based on their sexual orientation, which falls under the definition of hate speech.
LABEL: Hate
VARIANT: LLAMA DPO @ K=128
EXPLANATION: The post targets the homosexual community by using a derogatory slur "***" to demean and insult them. This language is offensive as it is rude, hurtful, and insulting towards individuals based on their sexual orientation. Therefore, the post is labeled as Offensive.
LABEL: Offensive
VARIANT: LLAMA +XMOD SFT+DPO @ K=256
EXPLANATION: The post targets the homosexual community by using a derogatory term "***" to insult and demean them. This language attacks and demeans a specific group based on their sexual orientation, which falls under the definition of hate speech. Therefore, the post is labeled as Hate.
LABEL: Hate
Figure 9: Examples of a post in HateXplain with gold label, along with the explanations of the T5 and Llama self-augmented variants at K=128, and the cross-trained model using all K=256 in total. We observe that models trained with more data (+Xmod) gives the correct classification.
POST:
it means people aren’t buying the lie! diversity k**s now prove me wrong
LABEL: IMPLICIT HATE
(a) A sample post in Latent Hate with gold label. Offensive content partially redacted.
VARIANT: T5 DPO @ K=128
EXPLANATION: This post does not involve any form of hatred, prejudice, or discrimination towards individuals or groups based on their characteristics. It is a statement expressing skepticism towards the belief that diversity ***, but it does not disparage or express any form of hate towards any specific person or group. Therefore, it can be classified as "Not Hate."
LABEL: Not Hate
VARIANT: T5 +XMOD SFT+DPO @ K= 256
EXPLANATION: This post does not involve any form of hatred, prejudice, or discrimination towards individuals or groups based on their characteristics. The post simply states that people are not buying the lie that diversity ***, and asks for proof to prove the opposite. There is no indication of any negative or discriminatory language towards any specific group or entity. Therefore, this post should be classified as "Not Hate."
LABEL: Not Hate
VARIANT: LLAMA DPO @ K=128
EXPLANATION: The post uses coded or indirect language to disparage a person or group based on their race. Overall, the use of the phrase "diversity ***" and the challenge to "prove me wrong" indicate a negative view of diversity, suggesting a belief that diversity is harmful. This coded language indirectly disparages people of color, making it an example of implicit hate.
LABEL: Implicit Hate
VARIANT: LLAMA +XMOD SFT+DPO @ K=256
EXPLANATION: This post should be classified as "Not Hate" because it does not involve any form of hatred, prejudice, or discrimination towards individuals or groups based on their characteristics. The post simply expresses a belief that diversity is a lie and challenges the reader to prove the speaker wrong.
LABEL: Not Hate
Figure 10: Examples of a post in Latent Hate with gold label, along with the explanations of the T5 and Llama self-augmented variants at K=128, and the cross-trained model using all K=256 in total. In this example, we observe that T5 absorbs the verbosity of Llama after cross-training, yet still gets the wrong label. On the other hand, Llama’s explanation is more terse after cross-training, and also classifies the post incorrectly.
POST: hi can i get a couple of grammes of your *** power ?
LABEL: INCITEMENT
(a) A sample post in Implicit Hate with gold label. Offensive content partially redacted.
VARIANT: T5 DPO @ K=128
EXPLANATION: The post targets minorities by expressing frustration over their perceived privilege and casting the majority group (*** people) as the real victims of racism. The post uses language that elevates *** power and unity, which is a known hate group ideology. The post also includes a sarcastic and ironic tone by asking for a "couple of grammes of your *** power," implying that minorities are inferior and need to be subjugated. This aligns with the definition of incitement, as it flaunts in-group unity and power while also elevating known hate groups and ideologies.
LABEL: Incitement
VARIANT: T5 +XMOD SFT+DPO @ K= 256
EXPLANATION: The post flaunts in-group unity and power by flaunting the idea of *** power. By using the term "*** power," the post elevates the idea of *** power and elevates the hate group ideology of *** supremacy. This aligns with the definition of incitement, which includes flaunting in-group unity and power or elevating known hate groups and ideologies.
LABEL: Incitement
VARIANT: LLAMA DPO @ K=128
EXPLANATION: This post targets *** people and their perceived desire for power and authority. By specifically asking for a couple of grammes of *** power, the post elevates and promotes the idea of *** supremacy, which is a hate group and ideology that advocates for the superiority of the *** race. This incitement of in-group unity and power aligns with the definition of incitement, as it promotes and supports a hate group and their extremist behavior.
LABEL: Incitement
VARIANT: LLAMA +XMOD SFT+DPO @ K=256
EXPLANATION: The post directly flaunts in-group unity and power by referencing *** power, which is a known hate group and ideology. The post specifically asks for a couple of grammes of *** power, which further promotes the idea of *** supremacy. This aligns with the definition of incitement, as it flaunts in-group unity and power and elevates known hate groups and ideologies.
LABEL: Incitement
Figure 11: Examples of a post in Implicit Hate with gold label, along with the explanations of the T5 and Llama self-augmented variants at K=128, and the cross-trained model using all K=256 in total. In this example, we also observe the cross-pollination of style in terms of verbosity, albeit to a lesser degree, but all models classify correctly.