### v3

# SMARTER: A Data‑efficient Framework to Improve Toxicity Detection with Explanation via Self-augmenting Large Language Models

Abstract
WARNING: This paper contains examples of offensive materials. To address toxic content on social media, we introduce SMARTER, a data-efficient 2-stage framework for explainable content moderation using Large Language Models (LLMs). In Stage 1, we leverage LLMs’ own outputs to generate synthetic explanations for correct and incorrect labels, enabling preference optimization with minimal supervision. In Stage 2, we refine explanation quality through cross-model training, allowing weaker models to align with stronger ones. Experiments on 3 classification tasks (HateXplain, Latent Hate, Implicit Hate) show SMARTER achieves up to 13% macro-F1 improvement over few-shot baselines using only 6-57% of training data. Our framework offers a scalable strategy for low-data settings by harnessing LLMs’ self-improvement for explainable moderation.
SMARTER: A Data‑efficient Framework to Improve Toxicity Detection with Explanation via Self-augmenting Large Language Models
Huy Nghiem, Advik Sachdeva, Hal Daumé III University of Maryland {nghiemh, asachde1, hal3}@umd.edu
1 Introduction
In recent years, social media platforms have enabled millions of users to virtually interact at a global scale. While a vital channel to disseminate information and knowledge, these platforms also facilitate the spread of harmful materials Hwang and Kim (2015); Cinelli et al. (2021); Nguyen et al. (2024). Arguably the most concerning among them is toxic content, which could induce serious negative psychological impact on the audience Castaño-Pulgarín et al. (2021); Windisch et al. (2022); Tontodimamma et al. (2021). As a result, content moderation is crucial to detect and mitigate its harm.
Toxic content encompasses a wide spectrum of terminologies whose definitions vary by platform: hate speech, cyberbully, sexist, racist etc. Gelber (2021); Anjum and Katarya (2024); Wang et al. (2025). Typically, human moderators must manually adjudicate the target content, an inefficient process taxing to the moderators themselves Baker et al. (2020); Spence et al. (2023). Recently, social media platforms can significantly reduce this overhead by integrating machine learning models to automate the detection pipeline d’Sa et al. (2020); Sumanth et al. (2022).
While capable of achieving impressive classification performance, these increasingly sophisticated models require ample resources to train, especially when the target label space covers nuanced concepts Fortuna et al. (2021); Sap et al. (2021); Zhao et al. (2021). Developing a robust dataset for training typically demands significant effort in curation and human annotation, often involving iterative updates to adapt the model to evolving linguistic trends Toraman et al. (2022); Bespalov et al. (2023); Rad et al. (2025). In addition, this pipeline is limited to decision-only pipeline. Failure to provide users with human-understandable signals renders the adjudication process opaque, inviting further issues on transparency and trustworthiness.
In this paper, we address the pressing need for intuitive explanations as a cornerstone of content moderation by leveraging Large Language Models (LLMs), which have shown remarkable capabilities in reasoning-related tasks Wei et al. (2021); Yao et al. (2024). Specifically, our contributions are:
  • $\diamond$
    We introduce SMARTER [^1] Our repository can be found at https://github.com/hnghiem-nlp/hate_dpo_public (Self-augMentAtion Regimen Towards Efficient Content ModeRation), a 2-stage framework where (1) each LLM self-augments on few-shot data using preference optimization with synthetic explanations, (2) cross-model refinement to further improvement by training one LLM with responses generated by another on held-out sets.
  • $\diamond$
    Experiments conducted on 3 classification tasks (HateXplain, Latent Hate, Implicit Hate) with Llama-3.1-8B-Instruct and COT-T5-XL demonstrate up to 13.5% increase in macro-F1 scores in few-shot settings with high-quality explanations, outperforming a range of commercial models.
While commercial APIs may incur high costs and limited transparency, SMARTER delivers competitive performance and explainability with 6-57% of training data, ensuring deployment control.
2 Related Work
Robust Toxicity Detection
Existing research Caselli et al. (2021); Sarkar et al. (2021); Toraman et al. (2022) has proposed strategies to enhance cross-domain robustness in toxicity detection, such as combining multiple datasets to improve generalization Antypas and Camacho-Collados (2023) and incorporating label definitions to support few-shot adaptation Nghiem et al. (2024). Building on these advances, recent studies leverage LLMs for toxicity detection, demonstrating strong zero- and few-shot performance Masud et al. (2024); Jahan et al. (2024); Roy et al. (2023); Kumarage et al. (2024). LLMs can also generate textual explanations, enhancing interpretability for moderators and users Di Bonaventura et al. (2024); Yang et al. (2023); Calabrese et al. (2024); Almohaimeed et al. (2025). We extend this line of work by training open-source LLMs to jointly perform classification and explanation to promote transparency.
Aligning LLMs to Human Preferences
In addition to extensive pretraining, modern LLMs also undergo post-training that both unlocks additional capabilities Longpre et al. (2023) and aligns the outputs towards human preferences Ouyang et al. (2022). Recent Reinforcement Learning (RL)-based techniques Shao et al. (2024); Zweiger et al. (2025); Nghiem et al. (2025) enable promising LLM self-alignment at expensive computational cost. Another body of work showcases LLMs to self-improve via their own synthetic data Tang et al. (2023); Li et al. (2023), such as Self-Refine Madaan et al. (2023) and Self-Instruct Wang et al. (2022). Building on these ideas, our framework presents a simple yet principled approach that allows LLMs to iteratively improve with minimal training cost while amenable to human supervision.
3 Overview of Framework
As shown in Algorithm 1, our framework consists of two stages. First, we focus on optimizing the LLMs’ classification performance through alignment tuning using self-augmenting synthetic data generation. In the second stage, we refine the quality of the generated explanations by cross-model refinement, leveraging the synergy between different LLMs. Section 4 and 5 present the experimental setup and findings for Stage 1. Section 6 focuses on cross-model improvement in Stage 2. Section 7 analyzes the consistency of explanations with respect to the definitions and labels.
Figure 1: Self-augmenting pipeline: for each post, explanations are conditionally generated for the gold label and all incorrect labels using prompt template in Figure 7. For DPO, data consists of matching the explanation of the correct label with another incorrect label’s.
Input: It means people aren’t buying the lie! diversity k**s now prove me wrong! Explanation: The post uses coded or indirect language to disparage a person or group based on their race. Overall, the use of the phrase "diversity k**s" and the challenge to "prove me wrong" indicate a negative view of diversity, suggesting a belief that diversity is harmful. This coded language indirectly disparages people of color, making it an example of implicit hate. Label: Implicit Hate
Figure 2: Example of explanation and classification label produced by Llama at $K=128$ shot on the dataset Latent Hate. For definition of categories, see Table 10.
Algorithm 1 High-Level Overview of SMARTER
1:Input: LLMs $\mathcal{L}$, performance metric $\mathcal{M}$ Datasets $\mathcal{D}_{\text{train}}\text{ (with explanation)},\mathcal{D}_{\text{val}},\mathcal{D}_{\text{test}}$,
2:Optional: Initial SFT pretraining on relevant data
3:Stage 1: Individual LLM Self-Augmentation
4:for $K\in\{K_{\text{start}},...,K_{\text{end}}$ } do
5:  Sample $K$-shot subset $\mathcal{D}^{(K)}$
6:  for all $\text{LLM}_{i}\in\mathcal{L}$ do
7:   SFT on $\mathcal{D}^{(K)}$
8:   Collect chosen responses for gold labels
9:   Collect rejected responses for incorrect labels
10:   Align via preference optimization using synthetic preference data; evaluate on $\mathcal{D}_{\text{val}}$
11:  end for
12:end for
13:Stage 2: Cross-Model Refinement
14:for $\text{LLM}_{a}\neq\text{LLM}_{b}\in\mathcal{L}$ do
15:  Select $K_{check}\in\{K_{\text{start}},...,K_{\text{end}}\}\rightarrow\mathcal{D}^{(K_{check})}$
16:  SFT $\text{LLM}_{a}$ on $\mathcal{D}^{(K_{check})}$
17:  Self-augment $\text{LLM}_{a}$ with responses from $\text{LLM}_{b}$ on $\mathcal{D}^{(K_{check}^{\prime})}\neq\mathcal{D}^{(K_{check})}$
18:end for
19:return Model with best performance $\mathcal{M}$ on $\mathcal{D}_{val/test}$
4 Preliminary Setup
In this section, we provide details about the empirical setup on data, model, and a description of classification technique via prompting.
4.1 Data
We select 2 datasets that cover distinct yet still relevant concepts to toxic content moderation and derive from them 3 classification tasks.
  • $\diamond$
    HateXplain, introduced by Mathew et al. (2021). Consisting of posts collected from 2 social media sites Gab and Twitter (now X), this dataset contains 20,148 samples annotated by crowdworkers of the Amazon Mechanical Turks (AMT) platform on 3 labels Hate, Offensive and Normal.
  • $\diamond$
    Latent Hate, curated by ElSherief et al. (2021), focuses on covert hate speech collected from Twitter. With 19,112 samples also annotated by AMT workers, this dataset’s primary label space contains 3 high-level categories: Not Hate, Implicit Hate, Explicit Hate. The authors also assigned the subset of 4,153 tweets of the Implicit Hate category a label among 6 secondary fine-grained categories: White Grievance, Incitement to Violence, Inferiority Language, Irony, Stereotypes and Misinformation, Threatening and Intimidation.
Label definitions are included, enabling context integration into explanations following Yang et al. (2023) and Nghiem and Daumé (2024). We denote these three tasks as HateXplain, Latent Hate, and Implicit Hate. Table 4 summarizes data splits; preprocessing details are in Appendix B. Figure 2 provides an example of the explanation produced by our model on Latent Hate.
4.2 Models
Extending our focus on accessibility, we employ 2 open-source LLMs that have exhibited strong performance in similar classification tasks:
  • $\diamond$
    COT-T5-XL: A variant of the base encoder-decoder Flan-T5-XL Longpre et al. (2023) architecture, this model consists of 3 billion parameters and is further trained on the COT corpus, a collection of 1.8 million samples with chain-of-thought style explanations, which enhances its capabilities on a variety of reasoning tasks Kim et al. (2023).
  • $\diamond$
    Llama-3.1-8B-Instruct: A decoder-only LLM released by Meta Dubey et al. (2024). Consisting of 8 billion parameters, this model is further instruction-tuned to aligned with human preferences in addition to being trained on extensive linguistic corpora.
For brevity, we refer to the models as T5 and Llama respectively for the remainder of this paper.
4.3 Classification via Prompting
We apply the Chain-of-Thought (COT) paradigm Wei et al. (2022) in our classification tasks to obtain the explanation from the LLMs necessary for content moderation. Inspired by Nghiem and Daumé (2024) and Yang et al. (2023), we use the prompt in Figure 7, which incorporates the target post, the set of labels corresponding to the task and their associated definitions as context for the LLM to provide its explanation and predicted category.
4.4 Auxiliary Pretraining
Nghiem and Daumé (2024) found that supervised finetuning (SFT) off-the-shelf LLMs on HateCOT – a general compilation of toxic posts, labels, definitions and synthetic explanations – can boost their performance in downstream hate speech detection. To maximize expected gains with minimal in-domain data, we adopt this strategy and perform SFT on the 2 chosen LLMs using HateCOT to prime them for our classification tasks. Finetuning is performed by training LoRA adapters Hu et al. (2021), a parameter-efficient approach implemented by the HuggingFace library (see Appendix F) as the basis for downstream experiments.
5 Stage 1: Individual Model Training
We demonstrate SMARTER’s data efficiency by training models at extreme low-shot settings and evaluating against off-the-shelf, commercial, and full-training baselines.
5.1 K-shot In-domain Finetuning
For each task, we select uniformly at random K-shot without replacement $\in\{16,32,64,128,256\}$ samples from the training set . We utilize the seed explanations by Nghiem and Daumé (2024) to further SFT the existing HateCOT-pretrained T5 and Llama models similar to the auxiliary pretraining phase above.
Self-augmenting via Alignment Tuning
We optimize the LLMs’ performance while not having access to more training data by exploiting the models’ sycophantic tendency. More concretely, we augment the training sample size by prompting the LLMs to generate explanations conditioned on the incorrect labels (as illustrated in Figure 1). This process supplies us with synthetic post-explanation pairs that are by design dispreferred compared to the true preferred sample pairs, enabling the use of the following alignment techniques to further finetune the models.
  • $\diamond$
    Direct Preference Optimization (DPO): An offline RL-based technique that optimizes the policy via an implicit reward model using preference data Rafailov et al. (2024). DPO requires human preference data where a pair of chosen vs. rejected responses is presented for each prompt.
  • $\diamond$
    Kahneman-Tversky Optimization (KTO): An alternative to DPO, KTO leverages prospect theory to construct human-aware losses (HALOs) to optimize the policy Ethayarajh et al. (2024). KTO only requires listwise human preference data: a data point consists of a prompt, a response and a binary flag that indicates the acceptability of the response.
Appendix F.3 offers more technical details on these techniques. Our self-augmenting methodology naturally procures training datasets in the required format for KTO by simply designating the explanation conditioned on the gold label as positive, and otherwise negative. For DPO, we match the explanations conditioned on the correct-incorrect pair sequentially to satisfy the pairwise chosen-rejected format. As an example, 100 posts of HateXplain (whose label space is of size 3), should produce 1003=300100*3=300 training samples for KTO, and 100(31)=200100*(3-1)=200 for DPO.
5.1.1 Experimental Pipeline
The experiments follow a two-stage process. We first finetune T5 and Llama at $K\in\{16,32,64,128\}$ followed by DPO self-augmentation. At $K=256$, we compare KTO and DPO alignment. [^2] KTO trainer for T5 is unavailable, so we evaluate KTO only for Llama. For comparison, we finetune both LLMs on full training data for classification without explanations (Full models), and similarly train ModernBERT (large) Warner et al. (2024) as an additional baseline.
In addition, we examine 2 partial sampling strategies to circumvent the proportionate growth of self-augmenting synthetic data to label space when K is large . In the first, we construct augmented data by selecting 128 and 192 shots of samples uniformly at random from the Baseline’s 256-shot pool. In the second, instead of $K$-shot sampling, we select $K^{\prime}*|S|*(|S|-1)$ samples from all possible post-label pairs of the 256-shot Baseline pool (incorrect labels included) to create augmented data, where $K^{\prime}\in\{128,192\}$ and $|S|$ is the label space’s size. While both approaches should yield identical size of training data for the same $K^{\prime}$, the first method, denoted DPO-K, preserves post-label parity whereas the second method, denoted DPO-N, yields higher diversity of post content.
Experiments are implemented with the HuggingFace library. We sample $K=50$ shot on the held out Validation extraction of each dataset to tune hyperparamters, and evaluate on samples from the Test set as shown in Table 4. Technical details are included in Appendix F.
5.2 Classification Results
5.2.1 Results for $K<=128$
We report macro-F1–unweighted mean of per-class F1 scores–as the evaluation metric Christen et al. (2023). Figure 3 displays the bar plots of the LLM’s Baseline and DPO-enhanced performance on the test sets.
Self-augmentation with DPO generally boosts classification performance over Baseline counterparts, with the sole reversal for T5 on Latent Hate at $K=16$ (2.1% reduction), along with static scores for T5 on HateXplain at $K=32$ and Llama on Latent Hate at $K=64$. These infrequent and negligible exceptions highlight our method’s consistent enhancement. For instance, F1 gains reach 14.5% for Llama (HateXplain, $K=128$) and 25% for T5 (Implicit Hate, $K=64$).
Notably, DPO self-augmentation improves performance even at $K\leq 64$, showcasing our method’s advantage in low-resource settings. Llama Baseline models appear to consistently outperform T5 across all tasks. However, with DPO self-augmentation, T5 is able to narrow the margin on Latent Hate and Implicit Hate. In contrast, Llama still dominates on HateXplain. These discrepancies suggest an interaction between dataset and model choice, and practitioners may benefit from contrasting different options with our self-augmenting techniques for best results.
Figure 3: Bar plots for $K$-shot classification experiments on 3 datasets using Llama and T5 models. Macro-F1 scores and percentage change over Baseline are displayed on top. Results for Baselines and DPO-augmented variants for $K\in\{16,32,64,128\}$ are displayed on the left subfigures. Results for $K=256$ of Baseline, KTO, DPO-augmented and its other variants on various sub-sampling strategies (section 5.1.1) are shown on the right. Horizontal lines show the F1 scores for Full models that use all training data.
5.2.2 Results for $K=256$
In Figure 3, we observe the macro-F1 scores for models $K=256$ on the subfigures on the right. DPO-augmentation continues to enhance Baseline models which already benefits from more training data relative to lower $K$ values, with improvement in scores reaching up to 13.2% for Llama and 14% for T5. In contrast, KTO-augmentation appears ineffectual on HateXplain and even hinders performance on Latent Hate and Implicit Hate, showing notable reduction from Baseline scores, respectively. With this, we hypothesize that DPO’s contrastive nature reinforces signals that allow the LLM to learn from its own mistakes, thereby sharpening its reasoning ability for this type of discriminatory task. On the other hand, KTO’s reliance on implicit signals from listwise data is not sufficient to improve discernment of classes when the semantic distinction between labels are nuanced for Latent Hate and Implicit Hate.
Results also show that DPO-augmented sub-sampling strategies still demonstrate generally better F1 scores than Baselines, further reinforcing the benefit of DPO-augmentation. There appears no clear winner between DPO-K and DPO-N among all 3 datasets. Augmenting with larger $K^{\prime}$ values (approaching 256) consistently yields scores closer to the maximal F1, indicating its role as a hyperparameter that practitioners need to decide for the trade-off between computational cost and benefit.
HateXplain Latent Hate Implicit Hate
Model F1 F1% Data% F1 F1% Data% F1 F1% Data%
SMARTER (Ours)
   Llama_DPO-256 0.64 89% 6% 0.69 100% 7% 0.60 90% 57%
   T5_DPO-256 0.62 86% 6% 0.65 94% 7% 0.59 88% 57%
Full Training Baselines
   Llama_Full 0.72 100% 100% 0.62 90% 100% 0.67 100% 100%
   T5_Full 0.72 100% 100% 0.61 88% 100% 0.64 96% 100%
   ModernBERT 0.70 97% 100% 0.61 88% 100% 0.64 96% 100%
Commercial: Zero-shot
   GPT-4o-mini 0.50 69% 0.54 78% 0.42 63%
   GPT-4.1 0.55 76% 0.60 87% 0.57 85%
   GPT-5-chat 0.56 78% 0.51 74% 0.58 87%
   Qwen-32B 0.54 75% 0.47 68% 0.49 73%
Commercial: 16-shot ICL
[…]
Ablation and Baseline Comparisons.
To isolate SMARTER’s contribution, Table 5 shows off-the-shelf models without HateCOT pretraining achieve 0.52 F1 on HateXplain. Adding HateCOT (Baseline $K=256$) improves to 0.58 (+0.06), while SMARTER’s DPO self-augmentation reaches 0.64 (+0.06 additional), contributing performance gain.
[…]
6 Stage 2: Cross-model Refinement
Having shown individual models self-improve efficiently, we explore whether cross-model training can further enhance performance by transferring explanation quality between LLMs. We first conduct human evaluation of both models’ explanations, then train each model on the other’s outputs.
Label T5 Llama Row Total
Normal 25 (25.5%) 73 (74.5%) 98
Offensive 63 (49.6%) 64 (50.4%) 127
Hate 54 (46.2%) 63 (53.8%) 117
Table 2: Annotator preferences for explanations on 342 HateXplain samples (majority vote).
6.1 Human Evaluation
Due to resource constraints, we only perform this evaluation pipeline for HateXplain. We opt to collect outputs from the SFT+DPO models at $K=128$ to preserve the rest of the training data for further refinement experiments. To make a fair comparison on the outputs to agree on the predicted label, we use the prompt in Figure 8 to obtain explanations conditioned on the correct gold label. This technique allows us to obtain 342 samples with consistent labels, or 89% of the possible 1283=384128*3=384 possible samples. The other 42 samples contain discrepant predicted labels, where the LLMs revert to behaviors ingrained by previous training.
To simulate real-life deployment, we solicit 14 crowdworkers from diverse demographic backgrounds on the platform Amazon Mechanical Turks (AMT) to annotate their preference on the explanations offered by the LLMs in this study . Annotators are encouraged to evaluate the pairs of explanation for each post based on the criteria of Clarity, Reasoning and Alignment as illustrated with the template in Figure 5. To avoid position bias, we present the pairs of explanations in randomized orders. All annotators are compensated fairly.
Annotation Results
Table 2 shows the annotation results, aggregated per sample by majority voting. Across the 3 labels in HateXplain, annotators show no significant preference for explanations by either model on Offensive and Hate. On the other hand, Llama’s explanations for Normal posts are overwhelmingly preferred.
6.2 Cross-model Refinement
Motivated by Llama’s superior human-rated explanations, we investigate whether models can benefit from each other’s outputs through cross-training. Below we describe our cross-model methodology, evaluate its impact on classification accuracy, and analyze how explanation styles evolve.
6.2.1 Cross-model Training Methodology
While Llama produces human-preferred explanations and is the natural deployment choice, we investigate whether models can adopt each other’s explanation styles without degrading classification performance. Using the prompting technique described in Section 6.1, we use the SFT+DPO $K=128$ variants of both Llama and T5 LLMs to generate explanations conditioned on the gold labels of the held-out 128-shot portion of the training data (unseen by either model). We then perform cross-model training: T5 is finetuned on Llama’s outputs and vice versa, applying both SFT and DPO self-augmentation as in Section 5.1.
6.2.2 Cross-model Classification Results
Figure 4: Macro-F1 scores on test portion of the 3 test sets for T5 and Llama cross-model refinement regimen. In each figure: the first and last bars are scores reported with only T5 model at $K=128$ and $K=256$ (directly from Figure 3); the middle bars are results after further finetuned using data from the complementary $K=128$ shots of counterpart model. Split-color bars: bottom color indicates the original model; top color indicates the counterpart model for additional cross-training.
Figure 4 shows the macro-F1 scores of both the original single-model SFT+DPO and cross-model finetuned variants on the test datasets.
Cross-model training enhances T5 while impairing Llama.
Applying SFT training using explanations from Llama improves performance over the baseline T5 SFT+DPO on HateXplain (F1 scores 0.62 from 0.60) and Latent Hate (0.64 from 0.62), with a slight reduction on Implicit Hate (0.54 from 0.56). On the other hand, applying SFT+DPO augmentation using Llama’s explanations from the complementary K=128 shot data allows T5 to exceed single-model performance on the full K=256 set across the board. The new maximum macro-F1 scores for T5 are 0.66 on HateXplain (up from 0.62 of single-model), 0.66 on Latent Hate (up from 0.65), and 0.61 on Implicit Hate (up from 0.59). Notably, these scores are superior to even Llama’s single-model performance at $K=256$ on HateXplain and Implicit Hate.
In contrast, SFT cross-training with T5’s output reduces Llama’s F1 scores to 0.58 from 0.63 on HateXplain while improving to 0.66 from 0.62 on Latent Hate and plateauing at 0.58 on Implicit Hate. Additional SFT+DPO cross-model augmentation slightly boosts performance on HateXplain and Implicit Hate while slightly impairing on Latent Hate. None of the cross-model Llama variants matches the single-models’ F1 scores at $K=256$.
6.2.3 Cross-model Style Analysis
Table 3: NLI-based consistency between explanations, predicted labels, and label definitions across datasets. Both DPO and cross-refined (XMOD) models show high entailment rates (>96%) with marginal percentages of contradictions (Contra.), neutral or undefined (Undef.) edge cases. Categories with all 0 values are omitted.
Label Consistency (%) Definition Consistency (%)
Dataset Model Train Entail Contra. Undef. Entail Contra. Neutral
HateXplain T5 DPO 99.2 0.8 0.0 98.2 1.5 0.3
XMOD 96.7 1.5 1.8 99.0 0.9 0.2
Llama DPO 97.5 1.2 1.3 97.2 2.6 0.2
XMOD 96.4 3.6 0.0 96.7 3.2 0.1
Latent Hate T5 DPO 98.8 1.2 0.0 97.6 2.4 0.0
XMOD 97.4 1.9 0.7 97.6 2.3 0.1
Llama DPO 97.3 2.7 0.0 97.3 2.3 0.3
XMOD 96.8 3.2 0.0 97.5 2.5 0.0
Implicit Hate T5 DPO 99.6 0.4 0.0 99.0 1.0 0.0
XMOD 98.3 1.4 0.4 97.5 2.0 0.5
Llama DPO 98.1 0.4 1.5 97.5 1.8 0.8
XMOD 96.9 2.8 0.1 96.0 3.6 0.4
We additionally train a style classifier based on the BERT architecture using explanations generated individually by T5 and Llama of the SFT+DPO variant at $K=128$ Devlin et al. (2019). A classification head is added to BERT, which leverages the [CLS] token from each explanation’s embedding to predict whether it was produced by T5 or LLaMA. Technical details are reported in Appendix F.3.
Figure 12 shows individually trained models exhibit their own styles, while cross-model SFT aligns outputs with the alternative’s style. Except for T5 on HateXplain, cross-model DPO slightly amplifies the base model’s style, suggesting this mechanism induces mild stylistic convergence.
Cross-model explanations may improve weak categories in weaker models.
We compare Precision, Recall, and F1-score of base and cross-trained models at $K=256$ (Figure 11) to assess cross-model refinement. Improvements often occur in categories where the paired model excels, indicating transferable reasoning patterns, while declines appear when the counterpart is weak. This asymmetry underscores our recommendation in line 19 of Algorithm 1 to validate and select the final model using a robust performance metric.
7 Analysis of Explanations
Using Natural Language Inference (NLI), we assess explanation consistency with labels and definitions, then analyze cross-model stylistic changes.
7.1 Analysis of Consistency
Explanation-Label Consistency
We prompt GPT-4o-mini with the template shown in Figure 9 to judge whether the offered explanations are logically consistent with the predicted label for each sample across 3 test sets via the 4 categories: Entail, Contradict, Neutral and Undefined (noise cases where either the predicted label or the explanation is missing). An author’s independent annotation of 150 random samples from this pool under the same instruction achieves nearly unanimous agreement with GPT-4o-mini (Cohen’s $\kappa=0.92$), an observation inline with recent works that validate LLM-as-a-judge approach in NLI applications Negru et al. (2025); Gallipoli and Cagliero (2025).
In Table 3, Entail values exceeding 96% indicate that explanations are largely consistent with their respective labels. However, cross-model training induces a marginal 2-3% increase in Contradiction compared to the DPO augmentation.
Explanation-Definition Consistency
Applying the same method to only the Entail samples from the previous analysis, using the prompt shown in Figure 10 (Cohen’s $\kappa=0.96$), we again observe consistently high entailment rates and a slight increase in contradiction under cross-model training that mirrors the aforementioned trend.
Overall, LLM-generated explanations remain predominantly consistent with both predicted labels and their definitions. Cross-model refinement can enhance classification but may introduce a slight risk of inconsistency, warranting periodic monitoring. Appendix E offers qualitative analysis and insights for further quality checks for deployment.
8 Discussion and Conclusion
Our empirical analysis yields several key recommendations to guide practitioners in deployment.
SMARTER delivers controllable, cost-efficient moderation.
Our staged pipeline achieves 86-100% of full-model performance using only 6-57% of training data (Table 1). Notably, we demonstrate this efficiency on challenging multi-class settings, whereas prior works tend to focus on simpler tasks Guo et al. (2023); Kim et al. (2024). Unlike commercial APIs—where 16-shot ICL exhibits high variance ($\pm$0.10) and can degrade performance (Table 6)—SMARTER produces deterministic, explainable outputs suitable for production deployment at a fraction of the cost. When labeled data is scarce and explainability is required, SMARTER provides cost-efficient, transparent production deployment versus commercial APIs.
Cross-model refinement enables targeted improvement.
T5’s superior cross-trained F1 scores (Figure 4) show that weaker models can adopt stronger models’ reasoning patterns. Practitioners facing category-specific weaknesses (Figure 11) can train multiple LLMs, then apply cross-model finetuning at high-quality checkpoints (e.g., $K=128$) to transfer strengths. Architectural diversity (encoder-decoder vs. decoder-only) appears beneficial, as Llama enhances reasoning diversity that T5 leverages Hao et al. (2024).
Human oversight remains essential.
Human evaluation (Table 2) shows category-specific preferences: Llama excels on Normal posts (3:1) while matching on Hate. Practitioners should periodically validate explanations, especially after cross-model training, to maintain consistency (Table 3).
We hope that SMARTER will enable practitioners to effectively and efficiently deploy explanation-based content moderation pipelines.
9 Limitations
We discuss several limitations in our research, along with their implications for future explorations.
Generalization beyond English
Our work focuses exclusively on English corpora. In reality, social media exists for every language. As both English-based LLMs and research literature often attract much more attention, we hope this framework will be adapted for other languages.
Augmentation to base framework
In our experiments, samples are randomly sampled from the label space. However, more strategic sampling choices (e.g., selecting based on how the level of classification difficulty calibrated on validation set) may boost performance. We encourage practitioners to explore augmentation to maximize the utility of our framework.
Model selection
While we perform extensive experiments, our selection of 2 models could be expanded. At the cost of more training resource, training more models with different architectures or generation offer extra options for performance comparison for mix-and-match cross-training. We encourage other researchers to experiment with more LLMs using our framework.
Limited scope of human validation
Due to budget constraints, we are only able to perform human validation for 1 dataset. In practice, it is recommended to solicit a larger pool of annotators to ensure appropriate amount of human oversight.
Risk of bias in automated explanation
While our framework aims to improve performance and versatility in content moderation, it does not alleviate the fundamental sensitive nature of this task. Content moderation could be misappropriated to suppress free speech or harm marginalized groups Dias Oliva (2020); Kozyreva et al. (2023). Other works have noted that LLMs exhibit bias behaviors in certain applications Nguyen et al. (2023); Gu et al. (2025). We urge researchers and practitioners to maintain vigilance when integrating our framework with human supervision to ensure ethical standards Lai et al. (2022); Cao et al. (2024).
10 Ethical Considerations
The authors are not aware of any ethical problems in the development of this work. This research uses publicly available dataset. We also further synonymize the content to the best of our ability to provide extra caution for privacy. We acknowledge that our framework and associated techniques could be abused for harmful purposes.
11 Acknowledgment
This work is funded by the NSF under Grant No. 2229885 (NSF Institute for Trustworthy AI in Law and Society, TRAILS). We thank the service of ACL ARR reviewers, area chairs and the editors of the ACL conference for our paper’s publication.
Appendix
Appendix A Amazon Mechanical Turk Annotation
We obtained IRB approval before soliciting annotation from the crowd workers. 14 workers were recruited to annotate the 342 explanations introduced in Section 6.1. Samples are assigned at random by the AMT platform to each annotator. Due to budget constraint, we opt to collect annotation for HateXplain to preserve statistical power instead of spreading across all 3 datasets. The breakdown of the annotators’ demographic is below:
  • $\diamond$
    Gender: Female (10), Male (3), Non-binary (1)
  • $\diamond$
    Age: 18-29 (2), 30-39 (5), 40-49 (5), 50+ (2)
  • $\diamond$
    Race/Ethnicity: Asian (3), Black (2), Hispanic/Latino (1), White (8)
  • $\diamond$
    Country of residence: United States (14)
  • $\diamond$
    Education level: 2-year college or equivalent (4), 4-year college or equivalent (6), High school or equivalent (3), Master degree or above (1)
Figure 5: Template to collect annotation for preference on explanations for Amazon Mechanical Turks crowdworkers.
Appendix B Data Pre-processing
The datasets used in this work are open-source. We peruse them as research artifacts accordingly to their license. To further mitigate risks of privacy, we anonymize posts by replacing user handles with the string <user> and remove URLs if they appear in the context of the posts.
We obtain the seed explanations from the repository associated with Nghiem and Daumé (2024) and merge them with the correct identifier keys.
Dataset Total Size Split Ratio K val K test Platform Label Space
HateXplain 20,148 60:20:20 50 400 G, T Normal, Offensive, Hate
Latent Hate 19,112 60:20:20 50 400 T Not Hate, Explicit Hate, Implicit Hate
Implicit Hate 4,153 50:20:30 50 150 T White Grievance, Incitement to Violence, Inferiority Language, Irony, Stereotypes and Misinformation, Threatening and Intimidation
Table 4: Datasets for evaluating our framework. Split Ratio designates the proportion of the train:validation:test split. K val and K test are the number of data points per class to perform $K$-shot sampling from the Validation and Test set. For Platform, G denotes Gab, T for X (formerly Twitter).
Appendix C Obtaining Explanations for In-domain Data
The prompt template in Figure 8 is used to collect explanation conditioned on a single label. Note that more labels (and their definitions) may also be added to this section as desired.
Dataset Model Macro F1
HateXplain Llama_OTS 0.52
T5_OTS 0.56
Latent Hate Llama_OTS 0.53
T5_OTS 0.42
Implicit Hate Llama_OTS 0.32
T5_OTS 0.33
Table 5: Comparison of off-the-shelf (OTS) model performance across datasets. LLMs are prompted to produce answers without explanations in this setting as shown in Figure 6.
### Instruction:
By considering their corresponding Definitions, label the following post with only one of these categories: {categories}.
Provide your response in the following format:
LABEL: [text]
### Definitions:
{definitions}
### Post:
{post}
### Response: [model output]
Figure 6: Prompt template for classification tasks without explanation generation. The model is instructed to directly output a label based on the provided definitions.
Dataset Model Min Mean Max Range
HateXplain GPT-4o-mini 0.22 0.29 0.42 0.20
GPT-4.1 0.47 0.52 0.60 0.13
GPT-5-chat 0.61 0.62 0.63 0.02
Qwen-32B 0.52 0.55 0.59 0.07
Latent Hate GPT-4o-mini 0.17 0.25 0.34 0.17
GPT-4.1 0.58 0.63 0.67 0.09
GPT-5-chat 0.54 0.60 0.65 0.11
Qwen-32B 0.54 0.57 0.60 0.06
Implicit Hate GPT-4o-mini 0.14 0.15 0.16 0.02
GPT-4.1 0.30 0.38 0.51 0.21
GPT-5-chat 0.29 0.40 0.51 0.22
Qwen-32B 0.44 0.48 0.51 0.07
Table 6: Detailed 16-shot in-context learning (ICL) results for commercial models across 3 random seeds. For each dataset, we report the minimum, mean, and maximum Macro F1 scores, along with the range (Max – Min) to illustrate performance variability. Models were prompted with 16 examples per class. High variance, particularly for GPT-4o-mini and on complex tasks like Implicit Hate, demonstrates ICL’s brittleness for nuanced multi-class toxicity detection.
### Instruction:
By considering their corresponding Definitions, label the following post with only one of these categories: categories. Provide a brief paragraph to explain step-by-step why the post should be classified with the provided Label based on the given Definitions.
If this post targets a group or entity relevant to the definition of the specified Label, explain who this target is/are and how that justifies the chosen Label.
Provide your message in the following format:
EXPLANATION: [text]
LABEL: [text]
### Definitions:
Hate: Language that targets [...]
Abusive: Language that is [...]
Normal: Does not contain [...]
### Post:
{post}
### Response: [model output]
Figure 7: Prompt template for classification tasks. The definition block should contain the set’s full label space
### Instruction:
By considering the following Post, the Label and its Definition, provide a brief paragraph to explain step-by-step why the post should be classified with the provided Label based on the given Definition.
If this post targets a group or entity relevant to the definition of the specified Label, explain who this target is/are and how that justifies the chosen Label.
Provide your message in the following format:
EXPLANATION: [text]
LABEL: [text]
### Definitions:
Hate: Language that targets [...]
### Post:
{post}
### Response:
Figure 8: Prompt template to obtain explanation conditioned on a single label and its definition.
### Instruction:
You are given a model’s prediction and its explanation for a social media post.
Your task is to judge whether the explanation is logically consistent with the predicted label.
Possible outputs:
- ENTAIL: The explanation supports and is faithful to the predicted label.
- CONTRADICT: The explanation contradicts the predicted label.
- NEUTRAL: The explanation is unrelated, vague, or does not clearly support the predicted label.
- UNDEFINED: Either the predicted label or the explanation is missing or empty.
Return only one word: ENTAIL, CONTRADICT, NEUTRAL, or UNDEFINED.
### Predicted Label:
{pred_label}
### Model Explanation:
{explanation}
### Response: [model output]
Figure 9: Prompt template for evaluating explanation–label consistency. The model is asked to judge whether a generated explanation is logically consistent with its predicted label.
### Instruction:
You are given a model’s predicted label and its corresponding explanation.
Your task is to judge whether the explanation is logically consistent with the definition of that label.
Possible outputs:
- ENTAIL: The explanation clearly aligns with and supports the definition.
- CONTRADICT: The explanation conflicts with the definition.
- NEUTRAL: The explanation is vague, unrelated, or does not clearly support the definition.
Return only one word: ENTAIL, CONTRADICT, or NEUTRAL.
### Official Definition:
{definition}
### Model Explanation:
{explanation}
### Response: [model output]
Figure 10: Prompt template for evaluating explanation–definition consistency. The model is asked to assess whether a generated explanation logically aligns with the formal definition of the predicted label.
Appendix D Stylistic Analysis
In Figure 12, we show the distribution of Llama vs T5 style pre- and post cross-model refined as recognized by the BERT style classifier.
(a) HateXlain
(b) Latent Hate
(c) Implicit Hate
Figure 11: Comparison of Precision, Recall, and Macro F1-scores for the T5 and Llama model families under two training regimes: self-augmentation only (Base) and cross-model refinement (+Xmod) at $K=256$. Models trained with cross-model refinement (right plot within each model family) generally show improvements in label categories where their counterparts already perform well, and conversely exhibit declines in areas corresponding to their weaknesses.
Appendix E Qualitative Analysis of Explanation
We examine the relatively small amount of non-Entail samples in Section 7 to observe general trends in the models’ inconsistency with respect to both the predicted label and definition.
First, T5 models typically exhibit higher consistency in explanation compared to their Llama counterparts as shown in Table 3. Even after cross-model refinement (XMOD), the percentage of Contradiction increases by at most 1% (Implicit Hate). While a conclusive explanation is out of scope, we hypothesize that T5’s encoder-decoder architecture supports higher robustness compared to Llama’s decoder-only.
Second, qualitative analysis reveals different patterns of inconsistency in the explanations offered by these 2 models. For HateXplain and Implicit Hate, Llama’s Contradict-tagged explanations tend to consistently mentions some negative and/or offensive sentiments in the post, yet ultimately yielding the conflicting label Normal (Figure 16). In contrast, T5 Contradict explanations – much fewer in overall number – generally asserts a label at the beginning yet explicitly rejects this label at the end. For the more challenging Implicit Hate, both models’ Contradict explanations typically contain "lost in context" inconsistency, where the justification is grounded in some other labels’ definition rather than the stated one, possibly due to the nuanced complexity of these finer-grained categories.
Our NLI-based consistency check also emerges as a simple yet useful quality check for downstream deployment.
In addition to classification metrics, moderators may opt to implement a secondary lightweight NLI-checker to alert when a threshold of inconsistency is reached.
Appendix F Technical Details
F.1 Hardware
All experiments are carried out on a maximum of 2 Nvidia RTX A6000 GPUs. We download the models from their Huggingface [^3] https://huggingface.co/ checkpoints. Finetuning is implemented with LoRA techniques Hu et al. (2021) via the TRL library  [^4] https://huggingface.co/docs/trl/en/index. The LoRA configurations for all settings are:
  • $\diamond$
    Rank: 64
  • $\diamond$
    alpha: 128
  • $\diamond$
    Dropout: 0.05
  • $\diamond$
    Target modules: $q$ and $v$ in projection layers
F.2 Inference Parameters
All models in our experiments use the following parameters during inference:
  • $\diamond$
    temperature: 0.0
  • $\diamond$
    max_tokens/max_new_tokens:
    • $\diamond$
      512 for classification with explanation
    • $\diamond$
      20 for classification without explanation
The library vLLM [^5] https://docs.vllm.ai/en/latest/ is used coupled with the openAI v1/chatcompletion endpoint to enable fast inference in our experiments. The OpenAI model GPT-5-chat-latest endpoint was accessed on September 12, 2025. We also use the Qwen2.5-32B-Instruct-AWQ version on HuggingFace.
F.3 Preference Optimization
Direct Preference Optimization
DPO is an alignment technique used to fine-tune large language models (LLMs) to better align with human preferences. Unlike traditional methods such as Reinforcement Learning from Human Feedback (RLHF), DPO simplifies the process by directly optimizing the policy model using a binary cross-entropy objective, thereby eliminating the need for a separate reward model and the complexities of reinforcement learning. This approach reparameterizes the reward model such that its optimal policy can be expressed in a closed form, transforming preference optimization into a classification problem Rafailov et al. (2024).
Kahneman-Tversky Optimization
KTO is another model alignment technique that aims to directly maximize the utility of a model’s generations by applying principles from Kahneman-Tversky prospect theory Kahneman and Tversky (2013). KTO evaluates outputs based on their perceived gains or losses relative to a reference point, incorporating the concept of loss aversion. This method utilizes data that simply indicates whether an output is desirable or undesirable for a given input, which can be more readily available and less expensive to collect than the paired preference data required by DPO Ethayarajh et al. (2024).
Implementation
We primarily use the implementation of the DPOTrainer and KTOTrainer classes via the TRL library hosted on the HuggingFace platform. For the DPOTrainer, we set $\beta=0.1$, and the loss to be default sigmoid loss. For the KTOTrainer, we similarly set the $\beta=0.1$ and use the default KTO loss option.
F.4 Hyperparameter Tuning
For ModernBERT, we use the HuggingFace Trainer and train for 4 epochs with AdamW optimizer. We use the associated tokenizer and pad the inputs to maximum length (512) for each batch.
For Base SFT finetuning, we use 3 epochs and learning rate 3e-4. We present the chosen parameters for DPO/KTO alignment tuning in Table 7, Table 8, Table 9 below.
T5 Llama
K Technique Epochs LR Epochs LR
16 DPO 3 5e-05 3 1e-05
32 DPO 3 5e-05 4 1e-05
64 DPO 3 1e-05 3 1e-05
128 DPO 3 5e-05 3 1e-05
256 DPO 4 1e-04 3 1e-05
256 KTO 3 5e-07
(a) Hyperparameters for regular training.
T5 Llama
K Technique Epochs LR Epochs LR
256 DPO-K128 3 1e-05 4 1e-04
256 DPO-K192 3 1e-04 1 7e-05
256 DPO-N128 3 5e-05 5 5e-06
256 DPO-N192 3 5e-05 5 5e-06
(b) Hyperparameters for subsampling training.
Table 7: HateXplain.
T5 Llama
K Technique Epochs LR Epochs LR
16 DPO 3 5e-05 3 1e-06
32 DPO 3 1e-04 3 1e-06
64 DPO 3 5e-05 3 1e-06
128 DPO 4 5e-05 3 1e-04
256 DPO 3 1e-04 3 1e-04
256 KTO 3 5e-07
(a) Hyperparameters for regular training.
T5 Llama
K Technique Epochs LR Epochs LR
256 DPO-K128 4 1e-04 4 1e-04
256 DPO-K192 3 1e-04 3 1e-04
256 DPO-N128 4 1e-04 3 1e-04
256 DPO-N192 4 1e-04 5 1e-04
(b) Hyperparameters for subsampling training.
Table 8: Latent Hate.
T5 Llama
K Technique Epochs LR Epochs LR
16 DPO 3 5e-05 3 5e-06
32 DPO 3 5e-05 4 5e-05
64 DPO 3 1e-04 3 1e-06
128 DPO 1 7e-05 3 1e-04
256 DPO 1 5e-05 1 5e-05
256 KTO 3 5e-07
(a) Hyperparameters for regular training.
T5 Llama
K Technique Epochs LR Epochs LR
256 DPO-K128 1 7e-05 1 1e-05
256 DPO-K192 1 7e-05 1 5e-05
256 DPO-N128 1 7e-05 1 1e-05
256 DPO-N192 1 7e-05 1 5e-05
(b) Hyperparameters for subsampling training.
Table 9: Implicit Hate
(a) HateXplain
(b) Latent Hate
(c) Implicit Hate
Figure 12: Percentage distribution of Llama- and T5-style explanations on test sets by BERT style classifiers, pre (Base) and post cross-model (Xmod) refined.
Dataset Definition
HateXplain $\diamond$ Normal: neither hate speech nor offensive. $\diamond$ Hate Speech: language which attacks, demeans, offends, threatens, or insults a group based on race, ethnic origin, religion, disability, gender, age, sexual orientation, or other traits. it is not the presence of certain words that makes the text hate speech, rather you should look the context the word is used in the text. $\diamond$ Offensive Language: usage of rude, hurtful, derogatory, obscene or insulting language to upset or embarrass people.
Latent Hate $\diamond$ Not Hate: speech or actions that do not involve any form of hatred, prejudice, or discrimination toward individuals or groups based on their characteristics. $\diamond$ Explicit Hate: openly expressed, direct forms of hatred and prejudice toward individuals or groups based on their characteristics. $\diamond$ Implicit Hate: coded or indirect language that disparages a person or group on the basis of protected characteristics like race, gender, and cultural identity.
Implicit Hate $\diamond$ White Grievance: includes frustration over a minority groups’ perceived privilege and casting majority groups as the real victims of racism. This language is linked to extremist behavior and support for violence. $\diamond$ Inferiority Language: implies one group or individual is inferior to another, and it can include dehumanization, denial of a person’s humanity, and toxic language that compares the target with disease, insects, animals. Related to assaults on human dignity, dominance, and declarations of superiority of the in group. $\diamond$ Irony: refers to the use of sarcasm , humor, and satire to attack or demean a protected class or individual. $\diamond$ Stereotypes And Misinformation: associate a protected class with negative attributes such as crime, or terrorism. includes misinformation that feeds stereotypes and vice versa, like Holocaust denial and other forms of historical negationism. $\diamond$ Threatening And Intimidation: conveys a speaker’s commitment to a target’s pain, injury, damage, loss or violation of rights, threats related to implicit violation of rights and freedoms, removal of opportunities, and more subtle forms of intimidation. $\diamond$ Incitement to Violence: includes flaunting in group unity and power or elevating known hate groups and ideologies
Table 10: Definitions of toxic label categories and definitions (collected from their original publications) across datasets.
Appendix G Examples of Explanations
We provide some examples of the explanations generated during our cross-training phase in Figure 13, Figure 14, and Figure 15.
POST:
what kind of brzen hapless *** uses an expression like "there’s no big *** there"
LABEL: HATE
(a) A sample post in HateXplain with gold label. Offensive content partially redacted.
VARIANT: T5 DPO @ K=128
EXPLANATION: The post targets the group of homosexuals by using a derogatory term "***" to insult and demean them. This language is offensive as it is rude, hurtful, and insulting. Therefore, the post is labeled as Offensive.
LABEL: Offensive
VARIANT: T5 +XMOD SFT+DPO @ K= 256
EXPLANATION: The post targets the group of individuals who identify as "brzen" and use the derogatory term "***" The language used in the post attacks and demeans this group based on their sexual orientation, which falls under the definition of hate speech.
LABEL: Hate
VARIANT: LLAMA DPO @ K=128
EXPLANATION: The post targets the homosexual community by using a derogatory slur "***" to demean and insult them. This language is offensive as it is rude, hurtful, and insulting towards individuals based on their sexual orientation. Therefore, the post is labeled as Offensive.
LABEL: Offensive
VARIANT: LLAMA +XMOD SFT+DPO @ K=256
EXPLANATION: The post targets the homosexual community by using a derogatory term "***" to insult and demean them. This language attacks and demeans a specific group based on their sexual orientation, which falls under the definition of hate speech. Therefore, the post is labeled as Hate.
LABEL: Hate
Figure 13: Examples of a post in HateXplain with gold label, along with the explanations of the T5 and Llama self-augmented variants at K=128, and the cross-trained model using all K=256 in total. We observe that models trained with more data (+Xmod) gives the correct classification.
POST:
it means people aren’t buying the lie! diversity k**s now prove me wrong
LABEL: IMPLICIT HATE
(a) A sample post in Latent Hate with gold label. Offensive content partially redacted.
VARIANT: T5 DPO @ K=128
EXPLANATION: This post does not involve any form of hatred, prejudice, or discrimination towards individuals or groups based on their characteristics. It is a statement expressing skepticism towards the belief that diversity ***, but it does not disparage or express any form of hate towards any specific person or group. Therefore, it can be classified as "Not Hate."
LABEL: Not Hate
VARIANT: T5 +XMOD SFT+DPO @ K= 256
EXPLANATION: This post does not involve any form of hatred, prejudice, or discrimination towards individuals or groups based on their characteristics. The post simply states that people are not buying the lie that diversity ***, and asks for proof to prove the opposite. There is no indication of any negative or discriminatory language towards any specific group or entity. Therefore, this post should be classified as "Not Hate."
LABEL: Not Hate
VARIANT: LLAMA DPO @ K=128
EXPLANATION: The post uses coded or indirect language to disparage a person or group based on their race. Overall, the use of the phrase "diversity ***" and the challenge to "prove me wrong" indicate a negative view of diversity, suggesting a belief that diversity is harmful. This coded language indirectly disparages people of color, making it an example of implicit hate.
LABEL: Implicit Hate
VARIANT: LLAMA +XMOD SFT+DPO @ K=256
EXPLANATION: This post should be classified as "Not Hate" because it does not involve any form of hatred, prejudice, or discrimination towards individuals or groups based on their characteristics. The post simply expresses a belief that diversity is a lie and challenges the reader to prove the speaker wrong.
LABEL: Not Hate
Figure 14: Examples of a post in Latent Hate with gold label, along with the explanations of the T5 and Llama self-augmented variants at K=128, and the cross-trained model using all K=256 in total. In this example, we observe that T5 absorbs the verbosity of Llama after cross-training, yet still gets the wrong label. On the other hand, Llama’s explanation is more terse after cross-training, and also classifies the post incorrectly.
POST: hi can i get a couple of grammes of your *** power ?
LABEL: INCITEMENT
(a) A sample post in Implicit Hate with gold label. Offensive content partially redacted.
VARIANT: T5 DPO @ K=128
EXPLANATION: The post targets minorities by expressing frustration over their perceived privilege and casting the majority group (*** people) as the real victims of racism. The post uses language that elevates *** power and unity, which is a known hate group ideology. The post also includes a sarcastic and ironic tone by asking for a "couple of grammes of your *** power," implying that minorities are inferior and need to be subjugated. This aligns with the definition of incitement, as it flaunts in-group unity and power while also elevating known hate groups and ideologies.
LABEL: Incitement
VARIANT: T5 +XMOD SFT+DPO @ K= 256
EXPLANATION: The post flaunts in-group unity and power by flaunting the idea of *** power. By using the term "*** power," the post elevates the idea of *** power and elevates the hate group ideology of *** supremacy. This aligns with the definition of incitement, which includes flaunting in-group unity and power or elevating known hate groups and ideologies.
LABEL: Incitement
VARIANT: LLAMA DPO @ K=128
EXPLANATION: This post targets *** people and their perceived desire for power and authority. By specifically asking for a couple of grammes of *** power, the post elevates and promotes the idea of *** supremacy, which is a hate group and ideology that advocates for the superiority of the *** race. This incitement of in-group unity and power aligns with the definition of incitement, as it promotes and supports a hate group and their extremist behavior.
LABEL: Incitement
VARIANT: LLAMA +XMOD SFT+DPO @ K=256
EXPLANATION: The post directly flaunts in-group unity and power by referencing *** power, which is a known hate group and ideology. The post specifically asks for a couple of grammes of *** power, which further promotes the idea of *** supremacy. This aligns with the definition of incitement, as it flaunts in-group unity and power and elevates known hate groups and ideologies.
LABEL: Incitement
Figure 15: Examples of a post in Implicit Hate with gold label, along with the explanations of the T5 and Llama self-augmented variants at K=128, and the cross-trained model using all K=256 in total. In this example, we also observe the cross-pollination of style in terms of verbosity, albeit to a lesser degree, but all models classify correctly.
LLAMA - HATEXPLAIN
EXPLANATION: The post targets the group of homosexuals by using the term "queer" in a derogatory manner. The language used attacks and demeans this group based on their sexual orientation, which falls under the definition of hate speech. Therefore, the post is labeled as "Normal."
LABEL: Normal
(a) Llama: The final label ‘Normal’ does not follow from the prior justification.
T5 - LATENT HATE
EXPLANATION: This post should be classified as "Not Hate" because it does not involve any form of hatred, prejudice, or discrimination towards individuals or groups based on their characteristics. The post simply mentions Alex Linder and white nationalism 1.0, without expressing any negative or discriminatory views towards any specific group or individual. Therefore, it does not meet the criteria for being classified as "Not Hate."
LABEL: Not Hate
(b) T5: The explanation starts with an asserted label yet later refutes this category.
LLAMA - IMPLICIT HATE
EXPLANATION: The post uses irony to attack and demean the idea of a wall, sarcastically suggesting that it should also prevent merpeople from the ocean. The use of humor and satire is used to mock the idea of a wall being a barrier to prevent all forms of transportation, including boats, tunnels, and airplanes, as well as the ocean.
LABEL: Threatening
(c) Llama: The explanation discusses factors directly relevant to ‘Irony’, yet the predicted label ‘Threatening’ is unrelated.
Figure 16: Examples of explanations that are judged to contains some forms of inconsistency with respect to the predicted label and/or its definition in our NLI-based analysis.